# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v3

@@ 2.2 DECaLS
Image processing was performed using Tractor (2016ascl.soft04008L) to extract sources. These sources are modelled using parametric profiles convolved with specific point spread functions (PSFs), including a delta function for point sources, an exponential profile, a de Vaucouleurs profile, and a Sérsic profile. The best-fit model magnitudes are used as the default photometry, with correction for Galactic extinction (1998ApJ...500..525S).
@@ 3.1 Overview of the previous PAC method
Suppose we aim to investigate two distinct populations of galaxies: one, denoted as ${\rm{pop}}_{1}$, is drawn from a spectroscopic catalogue, while the other, referred to as ${\rm{pop}}_{2}$, is sourced from a photometric catalogue. These populations are characterized within a relatively narrow redshift range. In our prior work 2022ApJ...925...31X, we introduced a method called PAC, which enables the accurate measurement of the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$, of objects with selected physical properties from ${\rm{pop}}_{2}$ around objects from ${\rm{pop}}_{1}$. It is based on
@@
For more details, see 2022ApJ...925...31X.
With this method, $\bar{n}_{2}w_{{\rm{p}}}$ is calculated in each redshift bin and then combined, potentially introducing two sources of error. First, the properties of ${\rm{pop}}_{2}$ are all computed using the mean redshift of the bin, while they are correlated with ${\rm{pop}}_{1}$ objects spanning the full width of the redshift bin. This can lead to inaccuracies in the properties due to differences in luminosity distances and the rest-frame wavelength covered by the photometric bands. Moreover, the measured $\omega_{12,\rm{w}}(\theta)$ is multiplied by the same $\bar{S}_{2}/r_{1}^{2}$ factor, even though it should vary for each source in ${\rm{pop}}_{1}$. Although, as demonstrated in 2022ApJ...925...31X, these approximations are found to be acceptable at high redshift, they become less accurate at lower redshifts ($z<0.1$) due to the much more significant fractional variation in $r_{1}$. Hence, we now present an improved PAC method that operates without the need for redshift bins.
We improve various aspects of the PAC method. To obtain accurate physical properties for ${\rm{pop}}_{2}$, we present a method to infer the physical properties for the entire photometric sample across the entire redshift range in Section 3.2. To eliminate the need for redshift bins and account for the variation in $\bar{S}_{2}/r_{1}^{2}$, we introduce a new estimator for the PAC method in Section 3.3. In Section 3.4, we employ colour cuts to exclude foreground and background galaxies from the photometric sample, enhancing the accuracy of our measurements. We use a new method in Section 3.5 to determine the stellar mass limits achievable with DECaLS, allowing us to fully exploit the available data. In Section 3.6, we assess the accuracy of the improved PAC method.
@@ 3.3 A new estimator for the PAC method
In the previous PAC method, for each redshift bin we first compute $\omega_{12,\rm{w}}$ using the Landy–Szalay estimator (1993ApJ...412...64L) with two corresponding random samples
@@
where $DD$, DRDR and $RR$ are the data-data, data-random and random-random pair counts. The angle $\theta$ varies for different pairs of $D_{1}$ and $R_{1}$ and is defined as $\theta=\arcsin{(r_{\rm{p}}/r_{1})}$. Subsequently, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$ is obtained using Equation 2 by multiplying it with $\bar{S}_{2}/r_{1}^{2}$, which is calculated using the mean redshift of each bin.
Instead, we now incorporate $\bar{S}_{2}$ and $r_{1}^{2}$ as weights of ${\rm{pop}}_{2}$ and ${\rm{pop}}_{1}$ sources. First, we calculate $\bar{S}_{2}$ for galaxies with varying properties at the saved redshifts. For instance, if we want to calculate $\bar{n}_{2}w_{{\rm{p}}}$ for ${\rm{pop}}_{2}$ within $N$ stellar mass bins at $z<0.2$ across 478 saved redshifts. This results in a total of $478\times N$ $\bar{S}_{2}$ measurements. Then, we compute $r_{1}^{2}$ for each source and random point of ${\rm{pop}}_{1}$. The weighted pairs are defined as:
@@ 3.5 Determining stellar mass limits
While it is evident that DECaLS can extend to fainter magnitude and lower mass limits, determining these values is essential for fully exploiting the dataset. However, this task is not straightforward as we are reaching unexplored territory, with limited knowledge about these faint galaxies. In our previous studies (2022ApJ...939..104X; 2023ApJ...944..200X), we first determined the $r$ or $z$ band galaxy depths of DECaLS. Then, we created apparent magnitude vs. stellar mass diagrams using GAMA DR4 data (2022MNRAS.513..439D) with $V_{\rm{max}}$ correction for galaxies at $z<0.2$ and using DES deep field data (2022MNRAS.509.3547H) with photoz for galaxies at $0.2<z<0.7$. Subsequently, we determined the stellar mass for which DECaLS achieves 95 per cent completeness at each redshift with the galaxy depths we obtained. These results are sensitive to the determination of galaxy depths, as well as to corrections for incompleteness, cosmic variance in GAMA, and photo-z errors in DES. Additionally, both GAMA and DESI struggle to measure enough galaxies with $M_{*}<10^{8.0}{\rm M}_{\sun}$, making this method impractical for the lower mass end.
@@
We repeat the process at various redshifts for $z<0.2$ for the entire, blue, and red populations. We present the results in the top panel of Figure 6 and also plot the results for the entire population from 2022ApJ...939..104X based on GAMA data for comparison. The results are discrete because we measure the angular surface density in stellar mass bins with an interval of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$. The stellar mass limits are nearly identical for the red and entire populations but are lower for blue galaxies. This is because red galaxies have higher mass-to-light ratios than blue ones, making them fainter at fixed stellar masses. As a result, the stellar mass limits for the entire population are primarily determined by the red galaxies. Compared to 2022ApJ...939..104X, we find that the stellar mass limits derived from GAMA are overly optimistic. This may suggest that GAMA does not fully sample the entire colour space of low-mass galaxies, potentially missing red ones. In the bottom panel, we further display the effective volume calculated using the maximum complete redshift, as well as with a redshift cutoff of $z>0.01$, for various stellar masses. Our results indicate the potential to investigate galaxies with stellar masses of $10^{6}{\rm M}_{\sun}$ within a sufficiently large volume using the PAC method with BGS and DECaLS.
@@
When utilizing low redshift data, it becomes necessary to re-evaluate the approximation in Equation 2. 2011ApJ...734...88W has demonstrated an accuracy of 3 per cent even at $r_{\rm{p}}=30h^{-1}\rm{Mpc}$ for various $r$-band absolute magnitude bins at $z=0.07$. We intend to expand the analysis to lower redshifts, encompass a broader range of galaxy populations, and incorporate additional sources of error that were not considered in 2011ApJ...734...88W.
The first error arises from the different definitions of $w_{{\rm{p}}}(r_{\rm{p}})$ and $\omega_{12}(\theta)$ that was comprehensively addressed in 2011ApJ...734...88W. Assuming we have a ${\rm{pop}}_{1}$ sample at a specific redshift, by definition, we have
@@
However, there is another source of error that has not been considered in 2011ApJ...734...88W. Suppose we are interested in obtaining measurements for ${\rm{pop}}_{2}$ within a specific stellar mass bin $M_{\rm{min}},M_{\rm{max}}]$ and the stellar masses of ${\rm{pop}}_{2}$ are calculated at the redshift of ${\rm{pop}}_{1}$ as described in Section 3.2. When we compute the angular correlation, ${\rm{pop}}_{1}$ is correlated with some galaxies with $M_{*}>M_{\rm{max}}$ at $r_{2}>r_{1}$ and $M_{*}<M_{\rm{min}}$ at $r_{2}<r_{1}$ because the masses of correlated galaxies are either underestimated or overestimated due to the inaccurate redshift assumed. This effect can impact the measurements because both $\bar{n}_{2}$ and $\xi(r_{12})$ depend on stellar mass, which is also negligible when $r_{1}$ is significantly larger than the correlation length. To investigate this effect, we re-write the $\omega_{12}(\theta)$ we measured as
@@
To estimate the accuracy of the approximation $\Delta\eta/\eta_{\rm{true}}$ for various redshifts and stellar mass bins, we employ the precise stellar-halo mass relation (SHMR) from 2023ApJ...944..200X and the CosmicGrowth $N$-body simulation (2019SCPMA..6219511J) with a subhalo catalogue constructed using HBT+ (2012MNRAS.427.2437H; 2016MNRAS.457.1208H). The CosmicGrowth simulation has a mass resolution of $5\times 10^{8}h^{-1}{\rm M}_{\sun}$ and a box size of $600\,h^{-1}\,\text{Mpc}$, which is appropriate for studying the distribution of galaxies with $M_{*}>10^{8}{\rm M}_{\sun}$ according to 2023ApJ...944..200X. We derive $\xi(r,M_{*})$ and $\Phi(M_{*})$ using the SHMR and simulation data, and then estimate $\Delta\eta/\eta_{\rm{true}}$ across different redshifts and for varying stellar masses of ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$. Figure 7 presents the results at 2 characteristic scales, $r_{\rm{p}}=1.25h^{-1}\rm{Mpc}$ and $12.5h^{-1}\rm{Mpc}$, across 5 redshifts $z=0.01,0.025,0.075$, and $0.1$, and within 18 stellar mass bins for ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$, spanning the range from $10^{8.0}{\rm M}_{\sun}$ to $10^{11.6}{\rm M}_{\sun}$. We find that $\Delta\eta/\eta_{\rm{true}}$ is smaller for smaller scales, lower ${\rm{pop}}_{2}$ mass, and higher redshift, with its dependence on the mass of ${\rm{pop}}_{1}$ being nearly negligible. There are indications that $\Delta\eta/\eta_{\rm{true}}$ begins to increase when $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$. This behaviour is attributed to the fact that within these mass bins, ${\rm{pop}}_{1}$ starts to correlate with some galaxies having $M_{\rm{pop}_{2}}<10^{8.0}{\rm M}_{\sun}$, whose host halos or subhalos are unresolved by the simulation. In general, we can maintain an accuracy of better than 3 per cent for almost all the stellar mass bins at $z=0.05$ within a scale of up to $12.5h^{-1}\rm{Mpc}$. It is worth noting that the measurements become less accurate in the most massive ${\rm{pop}}_{2}$ bins at lower redshifts. However, these redshifts are primarily intended for the study of low mass galaxies and have fewer massive galaxies. The PAC method remains highly accurate for $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$ even at $z=0.01$, and we anticipate that this level of accuracy extends to lower mass galaxies. Although the test above was conducted for the entire sample, we expect that the conclusion also holds for the red and blue sub-samples. This demonstrates the potential of the PAC method for exploring the unknown region of faint galaxies.
@@
Figure 10: Comparison of $w_{{\rm{p}}}$ at $8h^{-1}\rm{Mpc}$ for red $\rm{pop}_{2}^{\rm{s}}$} across various stellar mass bins. Each panel corresponds to one of the 11 $\rm{pop}_{1}$} stellar mass bins. The orange solid lines with shaded regions represent the results from the $w_{{\rm{p}}}$ of all $\rm{pop}_{2}^{\rm{s}}$} with $M_{*}<10^{10.8}{\rm M}_{\sun}$. For comparison, the projected matter correlations $w_{{\rm{p}}}^{\rm{mm}}$ at redshifts 0 (dashed lines) and 0.1 (dotted lines) calculated using CAMB (2000ApJ...538..473L) are also shown.
@@ 4 First Application with DESI Y1 BGS
In this section, we present the first application of the improved PAC method using the DESI Y1 BGS and DECaLS samples. We derive $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red ${\rm{pop}}_{2}$ galaxies down to stellar masses of $10^{5.3}{\rm M}_{\sun}$ and $10^{6.3}{\rm M}_{\sun}$, respectively. Following the approach of 2022ApJ...939..104X, we obtain measurements of the GSMFs for blue and red galaxies at these mass thresholds by combining $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ from the BGS dataset. We then combine these to obtain the total GSMF.
@@ 4.1 Methodology for measuring GSMF
2022ApJ...939..104X proposed a method to directly derive the galaxy GSMF by combining $\bar{n}_{2}w_{{\rm{p}}}$ from the PAC method with $w_{{\rm{p}}}$ from spectroscopic surveys, which may not be complete in stellar mass. In this approach, to match ${\rm{pop}}_{2}$ from the photometric sample, the corresponding ${\rm{pop}}_{2}^{\rm{s}}$ is selected from the spectroscopic sample within the same stellar mass bin and cross-correlated with ${\rm{pop}}_{1}$ to obtain $w_{{\rm{p}}}$. The GSMF is then estimated by comparing $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ within each stellar mass bin. This method is straightforward and relies on the key assumption that galaxy bias is primarily determined by stellar mass, such that ${\rm{pop}}_{2}^{\rm{s}}$ and ${\rm{pop}}_{2}$ have similar biases, even if ${\rm{pop}}_{2}^{\rm{s}}$ is highly incomplete. In this paper, we use a similar approach with several improvements and extensions.
Although $w_{{\rm{p}}}$ is primarily determined by stellar mass, residual scatter has been observed in numerous studies, which has been shown to depend on secondary properties such as colour and morphology (2006MNRAS.368...21L; 2016ApJ...819...58W; 2022ApJ...926..130X; 2022ApJ...928...10G). For example, massive red central galaxies are found in more massive halos compared to massive blue ones, resulting in stronger clustering (2016ApJ...819...58W; 2022ApJ...926..130X). Moreover, red galaxies have a higher satellite fraction than blue galaxies (2022ApJ...928...10G). These red satellites reside in more massive halos, which enhances the clustering of red galaxies at small scales (1-halo term) and slightly increases it at large scales (2-halo term). Consequently, at a fixed stellar mass, red galaxies exhibit higher clustering than blue galaxies, with the difference becoming more pronounced on smaller scales (2006MNRAS.368...21L). Therefore, the method used in 2022ApJ...939..104X may overestimate the GSMF, as red galaxies, with their higher mass-to-light ratios, are often among the first to be excluded when the sample becomes incomplete. This bias skews the sample toward a higher fraction of blue galaxies, leading to lower observed $w_{{\rm{p}}}$.
@@
Additionally, we also divide ${\rm{pop}}_{1}$ into different stellar mass bins following 2022ApJ...939..104X. For a given ${\rm{pop}}_{2}$ sample, the value of $\bar{n}_{2}$ from each ${\rm{pop}}_{1}$ sample should be consistent, as the method is independent of the properties of ${\rm{pop}}_{1}$. Therefore, comparing results from ${\rm{pop}}_{1}$ across different stellar mass bins serves as a robust test of our method. Additionally, refining ${\rm{pop}}_{1}$ into smaller stellar mass bins makes the $\bar{n}_{2}w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ measurements easier to interpret.

### v4

@@ 2.2 DECaLS
Image processing was performed using Tractor (Lang et al., 2016) to extract sources. These sources are modelled using parametric profiles convolved with specific point spread functions (PSFs), including a delta function for point sources, an exponential profile, a de Vaucouleurs profile, and a Sérsic profile. The best-fit model magnitudes are used as the default photometry, with correction for Galactic extinction (Schlegel et al., 1998).
@@ 3.1 Overview of the previous PAC method
Suppose we aim to investigate two distinct populations of galaxies: one, denoted as ${\rm{pop}}_{1}$, is drawn from a spectroscopic catalogue, while the other, referred to as ${\rm{pop}}_{2}$, is sourced from a photometric catalogue. These populations are characterized within a relatively narrow redshift range. In our prior work Paper I, we introduced a method called PAC, which enables the accurate measurement of the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$, of objects with selected physical properties from ${\rm{pop}}_{2}$ around objects from ${\rm{pop}}_{1}$. It is based on
@@
For more details, see Paper I.
With this method, $\bar{n}_{2}w_{{\rm{p}}}$ is calculated in each redshift bin and then combined, potentially introducing two sources of error. First, the properties of ${\rm{pop}}_{2}$ are all computed using the mean redshift of the bin, while they are correlated with ${\rm{pop}}_{1}$ objects spanning the full width of the redshift bin. This can lead to inaccuracies in the properties due to differences in luminosity distances and the rest-frame wavelength covered by the photometric bands. Moreover, the measured $\omega_{12,\rm{w}}(\theta)$ is multiplied by the same $\bar{S}_{2}/r_{1}^{2}$ factor, even though it should vary for each source in ${\rm{pop}}_{1}$. Although, as demonstrated in Paper I, these approximations are found to be acceptable at high redshift, they become less accurate at lower redshifts ($z<0.1$) due to the much more significant fractional variation in $r_{1}$. Hence, we now present an improved PAC method that operates without the need for redshift bins.
We improve various aspects of the PAC method. To obtain accurate physical properties for ${\rm{pop}}_{2}$, we present a method to infer the physical properties for the entire photometric sample across the entire redshift range in Section 3.2. To eliminate the need for redshift bins and account for the variation in $\bar{S}_{2}/r_{1}^{2}$, we introduce a new estimator for the PAC method in Section $3.$ In Section 3.4, we employ colour cuts to exclude foreground and background galaxies from the photometric sample, enhancing the accuracy of our measurements. We use a new method in Section 3.5 to determine the stellar mass limits achievable with DECaLS, allowing us to fully exploit the available data. In Section 3.6, we assess the accuracy of the improved PAC method.
@@ 3.3 A new estimator for the PAC method
In the previous PAC method, for each redshift bin we first compute $\omega_{12,\rm{w}}$ using the Landy–Szalay estimator (Landy and Szalay, 1993) with two corresponding random samples
@@
where DDDD, DRDR and RRRR are the data-data, data-random and random-random pair counts. The angle $\theta$ varies for different pairs of $D_{1}$ and $R_{1}$ and is defined as $\theta=\arcsin{(r_{\rm{p}}/r_{1})}$. Subsequently, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$ is obtained using Equation 2 by multiplying it with $\bar{S}_{2}/r_{1}^{2}$, which is calculated using the mean redshift of each bin.
Instead, we now incorporate $\bar{S}_{2}$ and $r_{1}^{2}$ as weights of ${\rm{pop}}_{2}$ and ${\rm{pop}}_{1}$ sources. First, we calculate $\bar{S}_{2}$ for galaxies with varying properties at the saved redshifts. For instance, if we want to calculate $\bar{n}_{2}w_{{\rm{p}}}$ for ${\rm{pop}}_{2}$ within NN stellar mass bins at $z<0.2$ across 478 saved redshifts. This results in a total of $478\times N$ $\bar{S}_{2}$ measurements. Then, we compute $r_{1}^{2}$ for each source and random point of ${\rm{pop}}_{1}$. The weighted pairs are defined as:
@@ 3.5 Determining stellar mass limits
[…]
@@
We repeat the process at various redshifts for $z<0.2$ for the entire, blue, and red populations. We present the results in the top panel of Figure 6 and also plot the results for the entire population from Paper III based on GAMA data for comparison. The results are discrete because we measure the angular surface density in stellar mass bins with an interval of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$. The stellar mass limits are nearly identical for the red and entire populations but are lower for blue galaxies. This is because red galaxies have higher mass-to-light ratios than blue ones, making them fainter at fixed stellar masses. As a result, the stellar mass limits for the entire population are primarily determined by the red galaxies. Compared to Paper III, we find that the stellar mass limits derived from GAMA are overly optimistic. This may suggest that GAMA does not fully sample the entire colour space of low-mass galaxies, potentially missing red ones. In the bottom panel, we further display the effective volume calculated using the maximum complete redshift, as well as with a redshift cutoff of $z>0.01$, for various stellar masses. Our results indicate the potential to investigate galaxies with stellar masses of $10^{6}{\rm M}_{\sun}$ within a sufficiently large volume using the PAC method with BGS and DECaLS.
@@
When utilizing low redshift data, it becomes necessary to re-evaluate the approximation in Equation 2. Wang et al. (2011) has demonstrated an accuracy of 3 per cent even at $r_{\rm{p}}=30h^{-1}\rm{Mpc}$ for various $r$-band absolute magnitude bins at $z=0.07$. We intend to expand the analysis to lower redshifts, encompass a broader range of galaxy populations, and incorporate additional sources of error that were not considered in Wang et al. (2011).
The first error arises from the different definitions of $w_{{\rm{p}}}(r_{\rm{p}})$ and $\omega_{12}(\theta)$ that was comprehensively addressed in Wang et al. (2011). Assuming we have a ${\rm{pop}}_{1}$ sample at a specific redshift, by definition, we have
@@
However, there is another source of error that has not been considered in Wang et al. (2011). Suppose we are interested in obtaining measurements for ${\rm{pop}}_{2}$ within a specific stellar mass bin $M_{\rm{min}},M_{\rm{max}}]$ and the stellar masses of ${\rm{pop}}_{2}$ are calculated at the redshift of ${\rm{pop}}_{1}$ as described in Section 3.2. When we compute the angular correlation, ${\rm{pop}}_{1}$ is correlated with some galaxies with $M_{*}>M_{\rm{max}}$ at $r_{2}>r_{1}$ and $M_{*}<M_{\rm{min}}$ at $r_{2}<r_{1}$ because the masses of correlated galaxies are either underestimated or overestimated due to the inaccurate redshift assumed. This effect can impact the measurements because both $\bar{n}_{2}$ and $\xi(r_{12})$ depend on stellar mass, which is also negligible when $r_{1}$ is significantly larger than the correlation length. To investigate this effect, we re-write the $\omega_{12}(\theta)$ we measured as
@@
To estimate the accuracy of the approximation $\Delta\eta/\eta_{\rm{true}}$ for various redshifts and stellar mass bins, we employ the precise stellar-halo mass relation (SHMR) from Paper IV and the CosmicGrowth NN-body simulation (Jing, 2019) with a subhalo catalogue constructed using HBT+ (Han et al., 2012, 2016). The CosmicGrowth simulation has a mass resolution of $5\times 10^{8}h^{-1}{\rm M}_{\sun}$ and a box size of $600\,h^{-1}\,\text{Mpc}$, which is appropriate for studying the distribution of galaxies with $M_{*}>10^{8}{\rm M}_{\sun}$ according to Paper IV. We derive $\xi(r,M_{*})$ and $\Phi(M_{*})$ using the SHMR and simulation data, and then estimate $\Delta\eta/\eta_{\rm{true}}$ across different redshifts and for varying stellar masses of ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$. Figure 7 presents the results at 2 characteristic scales, $r_{\rm{p}}=1.25h^{-1}\rm{Mpc}$ and $12.5h^{-1}\rm{Mpc}$, across 5 redshifts $z=0.01,0.025,0.075$, and $0.1$, and within 18 stellar mass bins for ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$, spanning the range from $10^{8.0}{\rm M}_{\sun}$ to $10^{11.6}{\rm M}_{\sun}$. We find that $\Delta\eta/\eta_{\rm{true}}$ is smaller for smaller scales, lower ${\rm{pop}}_{2}$ mass, and higher redshift, with its dependence on the mass of ${\rm{pop}}_{1}$ being nearly negligible. There are indications that $\Delta\eta/\eta_{\rm{true}}$ begins to increase when $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$. This behaviour is attributed to the fact that within these mass bins, ${\rm{pop}}_{1}$ starts to correlate with some galaxies having $M_{\rm{pop}_{2}}<10^{8.0}{\rm M}_{\sun}$, whose host halos or subhalos are unresolved by the simulation. In general, we can maintain an accuracy of better than 3 per cent for almost all the stellar mass bins at $z=0.05$ within a scale of up to $12.5h^{-1}\rm{Mpc}$. It is worth noting that the measurements become less accurate in the most massive ${\rm{pop}}_{2}$ bins at lower redshifts. However, these redshifts are primarily intended for the study of low mass galaxies and have fewer massive galaxies. The PAC method remains highly accurate for $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$ even at $z=0.01$, and we anticipate that this level of accuracy extends to lower mass galaxies. Although the test above was conducted for the entire sample, we expect that the conclusion also holds for the red and blue sub-samples. This demonstrates the potential of the PAC method for exploring the unknown region of faint galaxies.
@@
Figure 10: Comparison of $w_{{\rm{p}}}$ at $8h^{-1}\rm{Mpc}$ for red $\rm{pop}_{2}^{\rm{s}}$} across various stellar mass bins. Each panel corresponds to one of the 11 $\rm{pop}_{1}$} stellar mass bins. The orange solid lines with shaded regions represent the results from the $w_{{\rm{p}}}$ of all $\rm{pop}_{2}^{\rm{s}}$} with $M_{*}<10^{10.8}{\rm M}_{\sun}$. For comparison, the projected matter correlations $w_{{\rm{p}}}^{\rm{mm}}$ at redshifts 0 (dashed lines) and 0.1 (dotted lines) calculated using CAMB (Lewis et al., 2000) are also shown.
@@ 4 First Application with DESI Y1 BGS
In this section, we present the first application of the improved PAC method using the DESI Y1 BGS and DECaLS samples. We derive $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red ${\rm{pop}}_{2}$ galaxies down to stellar masses of $10^{5.3}{\rm M}_{\sun}$ and $10^{6.3}{\rm M}_{\sun}$, respectively. Following the approach of Paper III, we obtain measurements of the GSMFs for blue and red galaxies at these mass thresholds by combining $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ from the BGS dataset. We then combine these to obtain the total GSMF.
@@ 4.1 Methodology for measuring GSMF
Paper III proposed a method to directly derive the galaxy GSMF by combining $\bar{n}_{2}w_{{\rm{p}}}$ from the PAC method with $w_{{\rm{p}}}$ from spectroscopic surveys, which may not be complete in stellar mass. In this approach, to match ${\rm{pop}}_{2}$ from the photometric sample, the corresponding ${\rm{pop}}_{2}^{\rm{s}}$ is selected from the spectroscopic sample within the same stellar mass bin and cross-correlated with ${\rm{pop}}_{1}$ to obtain $w_{{\rm{p}}}$. The GSMF is then estimated by comparing $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ within each stellar mass bin. This method is straightforward and relies on the key assumption that galaxy bias is primarily determined by stellar mass, such that ${\rm{pop}}_{2}^{\rm{s}}$ and ${\rm{pop}}_{2}$ have similar biases, even if ${\rm{pop}}_{2}^{\rm{s}}$ is highly incomplete. In this paper, we use a similar approach with several improvements and extensions.
Although $w_{{\rm{p}}}$ is primarily determined by stellar mass, residual scatter has been observed in numerous studies, which has been shown to depend on secondary properties such as colour and morphology (Li et al., 2006; Wang et al., 2016b; Xu and Jing, 2022; Gao et al., 2022). For example, massive red central galaxies are found in more massive halos compared to massive blue ones, resulting in stronger clustering (Wang et al., 2016b; Xu and Jing, 2022). Moreover, red galaxies have a higher satellite fraction than blue galaxies (Gao et al., 2022). These red satellites reside in more massive halos, which enhances the clustering of red galaxies at small scales (1-halo term) and slightly increases it at large scales (2-halo term). Consequently, at a fixed stellar mass, red galaxies exhibit higher clustering than blue galaxies, with the difference becoming more pronounced on smaller scales (Li et al., 2006). Therefore, the method used in Paper III may overestimate the GSMF, as red galaxies, with their higher mass-to-light ratios, are often among the first to be excluded when the sample becomes incomplete. This bias skews the sample toward a higher fraction of blue galaxies, leading to lower observed $w_{{\rm{p}}}$.
@@
Additionally, we also divide ${\rm{pop}}_{1}$ into different stellar mass bins following Paper III. For a given ${\rm{pop}}_{2}$ sample, the value of $\bar{n}_{2}$ from each ${\rm{pop}}_{1}$ sample should be consistent, as the method is independent of the properties of ${\rm{pop}}_{1}$. Therefore, comparing results from ${\rm{pop}}_{1}$ across different stellar mass bins serves as a robust test of our method. Additionally, refining ${\rm{pop}}_{1}$ into smaller stellar mass bins makes the $\bar{n}_{2}w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ measurements easier to interpret.